Calculate the resonance frequency of a circuit capacitive reactance value of a 310nF capacitor and inductive reactance value of a 2m Ω inductance.


1
Expert's answer
2021-04-26T18:45:13-0400

"C=310nf=310\times 10^{-9}F,



L=2m\Omega=2\times 10^{-3} \Omega"


Resonance frequency "\omega_o=\dfrac{1}{\sqrt{LC}}=\dfrac{1}{\sqrt{310\times 2\times 10^{-12}}}=\dfrac{10^{12}}{\sqrt{620}}=1.62\times 10^9 sec^{-1}"